The apparent magnitude of star at distance of 100 parsec is 7. what is its absolute magnitude?
1
Expert's answer
2012-11-09T04:00:48-0500
The absolute magnitude is magnitude the object would have ifit was at distance of 10 parsecs.
This is 10 times closer then object is. At this distance it would have 100
times larger luminocity. The difference in magniеudes can be find as
m1-m2 = -2.5 lg(L1/L2)
Hence, its absolute magnitude is
m1 = 7-2.5 ln(100) = 7-5 = 2